OpenStudy (anonymous):

what is the solution of the system: x-2y=2 ; 4x-8y=8

OpenStudy (uri):

Substitution?

OpenStudy (anonymous):

i have to graph it, so substitution or elimination

OpenStudy (anonymous):

Notice that equation #2 is just equation #1 with both sides multiplied by 4. So, what you actually have here is the same line with 2 representations. You can think of one line being placed on top of the other. This is a dependent system with an infinite number of solutions. To graph it, choose either line and graph that line.

OpenStudy (anonymous):

Wow, I didn't catch that. But I still have to put the equation in y=mx+b form, correct?

OpenStudy (anonymous):

That is the simplest way to graph. It's the way I would do it. Another way is to plot both the x and y intercepts and draw a line through them.

OpenStudy (anonymous):

(4,0) (0,-8) ?

OpenStudy (anonymous):

No. Take either equation and stick with just the one equation. First set one variable to "0", plug in that "0" where you see that variable and solve for the other variable. Then repeat the process with the other variable.

OpenStudy (anonymous):

ohhhh okay, (2,-1)

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

No, if you show your work, I'll be able to see what you are doing wrong.

OpenStudy (anonymous):

crap....okay i did: 4(0)-8y=8 -8y=8 y=-1 4x-8(0)=8 4x=8 x=2

OpenStudy (anonymous):

What you did above is correct, except you represented the intercepts wrong. Once you get those 2 values, the intercepts aren't: (2, -1) They are: (2, 0) and (0, -1) two distinct points.

OpenStudy (anonymous):

omg I always do that!! ugh, thank you!!

OpenStudy (anonymous):

uw!